• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name Olivier

    Log entry time 23:17:59 on July26,2003

    Entry number 107198

    keyword=BCM calibration delayed

    The BCM calibration that was supposed to be taken on sunday owl shift will not happen tonight, due to the delay accumulated at the beginning of the 4 pass running. If things go smoothly tomorrow (as they have been in the past 12 hours or so), we might be able to complete the physics program at 4 pass during monday owl. In this case, we MAY do the BCM calibration at the end of monday owl.

    Otherwise, it is postponed until later.